QUESTION 47

  1. Determine the equilibrium constant for the following reaction at 581 K.

    HCN(g) + 2 H2(g) → CH3NH2(g)      ΔH° =  -158 kJ;  ΔS°= -219.9 J/K

    3.99 × 1012

    3.26 × 10-12

    13.0

    2.97 × 1011

    0.002

    4.92 × 1025

    3.4 × 10-12

    None of these

    523.2

QUESTION 48

  1. The standard emf for the cell using the overall cell reaction below is +2.20 V:

                2Al(s) + 3I2 (s)  →  2Al3+ (aq) + 6I- (aq)

    The emf generated by the cell when [Al3+] = 1.3 × 10-4 M and [I-] = 0.044 M is ________ V.

    2.18

    2.20

    2.25

    none of these

    2.32

    2.51

    2.36

    2.67
Answer #1

047 HENLg + 24 g → CH₃NH2 AH = -15 863) AS = -219. 3 Jk AG-AH-TAS → Gibbs Brelo energy AGO -15B X183 – (581) (-219,9) 3o -30

444 EU = +2:24 6 Ecell = 2.2 + 0.059 (15.911) 2Al(s) + 34 6 DAH b) +6 I lag. CAPA] = 13810-4 m C1-1=0.044 M Exay__2:356 V Eco